# Joint estimation of a two-phase spin rotation beyond classical limit

### v1

@@ Abstract
Quantum metrology employs entanglement to enhance measurement precision [1, 2, 3]. The focus and progress so far have primarily centered on estimating a single parameter. In diverse application scenarios, the estimation of more than one single parameter is often required. Joint estimation of multiple parameters can benefit from additional advantages for further enhanced precision. Here we report quantum-enhanced measurement of simultaneous spin rotations around two orthogonal axes, making use of spin-nematic squeezing in an atomic Bose-Einstein condensate. Aided by the $F=2$ atomic ground hyperfine manifold coupled to the nematic-squeezed $F=1$ states as an auxiliary field through a sequence of microwave (MW) pulses, simultaneous measurement of multiple spin-1 observables is demonstrated, reaching an enhancement of 3.3 to 6.3 decibels (dB) beyond the classical limit over a wide range of rotation angles. Our work realizes the first enhanced multi-parameter estimation using entangled massive particles as a probe. The techniques developed and the protocols implemented also highlight the application of two-mode squeezed vacuum states in quantum-enhanced sensing of noncommuting spin rotations simultaneously.
@@
Quantum-enhanced metrology improves parameter estimation precision in sensing beyond the standard quantum limit (SQL), or the classical limit [1, 2, 3]. The fruitful recent achievement at advanced LIGO with squeezed light [4], yielding an increased number of events detected [5], has made quantum-enhanced metrology one of the most promising quantum technologies in the near future. It has been demonstrated in experiments with photons [6], neutral atoms [7, 8], trapped ions [9, 10] and artificial atoms [11]. Implemented at high-sensitivity atom interferometers, such as atomic clocks [12] and atomic magnetometers [13], significant technological gains will likely occur soon.
@@
Ideas for quantum-enhanced sensing have been extended to multi-parameter settings as well, where estimations of multiple parameters can simultaneously reach beyond SQL with suitable quantum entanglement [14, 15]. These extensions are relevant for a broad range of applications, such as imaging and microscopy [16, 17], multi-dimensional field sensing [18, 19], and quantum networks for atomic clocks [20] etc., and they have motivated growing interests from both theoretical and experimental perspectives. A significant challenge in multi-parameter estimation arises when the generators of parameter-encoding transformations do not commute. Observables that provide maximal information about different parameters need not necessarily commute [21], which could potentially impair quantum-enhanced sensitivities for all the parameters. Recently, Lorcán O. Conlon, et al. have experimentally demonstrated optimal single- and two-copy collective measurements for simultaneously estimating two noncommuting qubit rotations [22]. On the other hand, a two-mode Einstein-Podolsky-Rosen (EPR) entangled system exhibits perfect correlations not only between the positions but also between their momenta. For example, the difference $x_{1}-x_{2}$ of the position operators $x_{1}$ and $x_{2}$ commutes with the sum $p_{1}+p_{2}$ of the momentum operators $p_{1}$ and $p_{2}$, likewise for $x_{1}+x_{2}$ and $p_{1}-p_{2}$. Such EPR correlation is known to be capable of exhibiting squeezed noise in two conjugate observables, which allows for the simultaneous estimation of two encoded parameters beyond the classical limit. The original idea was first delineated in the form of quantum dense coding [23] and was later demonstrated experimentally with optical parametric amplifiers [24, 25, 26, 27]. A related recent study [28] reports estimating displacements beyond the SQL along two quadratures for the center-of-mass motion of a trapped ion. We also notice that a two-mode squeezed state can be also prepared with trapped ion system with a laser-free method using additional radio potentials, which is utilized to demonstrate the SU(1,1) interferometers [29].
@@
Figure 1: Spin-nematic squeezing in both SU(2) subspaces. (a) The generation of spin-nematic squeezed state is illustrated by quasi-probability distributions on both SU(2) subspace Bloch spheres, respectively spanned by $\left\{L_{y},Q_{xz},\left(Q_{xx}-Q_{zz}\right)/2\right\}$ and $\left\{Q_{yz},L_{x},\left(Q_{yy}-Q_{zz}\right)/2\right\}$, simulated here for ${N=100}$ atoms. The initial polar state corresponds to all atoms in the $m_{F}=0$ component. After a short evolution time at $q\approx|c_{2}|$, spin-nematic squeezing develops along the separatrix (dark gray lines) on both Bloch spheres. (b) The measured squeezing parameter at the optimal squeezing (for $q=|c_{2}|$) versus evolution time. The gray dots show the experimental results of measured squeezing in a single SU(2) subspace by applying an RF rotation to couple sublevels in the $F=1$ manifold [31]. The blue and orange dots show the experimental results of simultaneously measured squeezing in both SU(2) subspaces by the MW pulse sequence using the $F=2$ manifold as an ancillary field. The squeezing parameter of the simultaneous measurement scheme diminishes at longer times due to imperfections of the MW pulses as well as other decoherence or dissipation mechanisms. All markers denote experimental results averaged over 100 runs and the corresponding solid and dashed lines denote numerical calculations based on truncated Wigner approximation. (c) The lower panel shows the MW pulse measurement sequence. The complete measurement scheme lasts for about 374 $\mu$s in a single experiment run. The blue, yellow, and green squares correspond to the MW pulses with $\sigma_{+}$,$\sigma_{-}$, and $\pi-$polarization, respectively. The upper panel illustrates atom number distributions in all sublevels before (empty shape) and after (filled shape) the MW pulse is implemented. The blue and red shapes denote the signals and pump fields for atomic homodyne detection, while the purple shapes depict the atoms after $\pi/2$-pulse couplings between the signals and the pumps. A small number of atoms remain in the $|F=1,m_{F}=0\rangle$ state due to due to imperfections of the pulse sequences. Here $C^{i,j}_{\varPhi}$ denotes the coherent coupling between $\left|F=1,m_{F}=i\right\rangle$ and $\left|F=2,m_{F}=j\right\rangle$ at a rotation angle $\varPhi$. For the first three pulses, $\varPhi_{1}$,$\varPhi_{2}$, and $\varPhi_{3}$ are adjusted by changing the power ratio between $\sigma_{+}$ and $\sigma_{-}$ polarizations to approximately split the atoms in $\left|F=1,m_{F}=0\right\rangle$ into two halves coherently.
@@
The quantum-enhanced joint estimation we report concerns a spin rotation characterized by two unknown phases encoded into noncommuting collective spin components $L_{x}$ and $L_{y}$ of a spinor atomic Bose-Einstein condensate (BEC). The EPR-like entanglement is revealed to be inherent in the spin-nematic squeezed state generated by atomic spin-mixing [30, 31, 32], which is found to exhibit squeezed quantum noise in both spin-nematic SU(2) subspaces [33]. The associated correlation is demonstrated experimentally capable of beating the SQL precision by joint estimation of the two rotation phases simultaneously in this work. To read out the phases encoded into the probe state, we implement a protocol to measure the two noncommuting rank-2 quadrupole moments $Q_{xz}$ and $Q_{yz}$ of the collective spin observables simultaneously, reminiscent of atomic homodyne detection [30], by taking the $F=2$ hyperfine states as the auxiliary ones and applying a sequence of microwave pulses to transform $Q_{xz}$ and $Q_{yz}$ into the population differences of atomic hyperfine sublevels while preserving the squeezed noise correlation. With the protocol to be detailed later, both $Q_{xz}$ and $Q_{yz}$ are measured in a single experimental run, and the metrological gains for joint estimation of the two phases reach $3.3$ dB to $6.3$ dB beyond the SQL over a wide range of rotations explored, where the SQL $1/\sqrt{2N}$ is for estimating two parameters with an ensemble of $N$ spin-1 particles [34] (see Supplementary Material). Consistent with the standard practice of estimating a single parameter beyond the classical limit, we carry out strong (collapsing) measurements to obtain the maximal amount of information by adopting an atomic homodyne protocol. In Ref. [35], the back actions on the state for weak continuous measurement of noncommuting observables are studied.
@@
Figure 2: Joint estimation of spin rotation angles around $\hat{L}_{x}$ and $\hat{L}_{y}$. (a) Schematic of the joint estimation interferometry protocol with quasiprobability distributions for the respective states plotted in the $\left\{L_{x},Q_{yz}\right\}$ and $\left\{L_{y},Q_{xz}\right\}$ planes. $U_{\mathrm{S}}$, $U_{\mathrm{R}}$, and $U_{\vec{\varphi}}$ denote the unitary evolutions generated by spin-mixing, spinor phase rotation, and encoding spin rotations, respectively. (b) The distributions of the simultaneously measured data for the polar state (red dots) and spin-nematic squeezed state (blue dots) sampled over 100 times after spin rotations at different $\varphi_{1}$ and $\varphi_{2}$ encoded. The distribution of the polar state is much broader than that of the spin-nematic squeezed state, hence the higher sensitivities for the estimation from the latter. Here $\varphi_{1}=\alpha\cos\Theta$ and $\varphi_{2}=\alpha\sin\Theta$. The red (blue) circles delineate the boundary of the measured data, $Q_{yz}^{\mathrm{meas}}$ and $Q_{xz}^{\mathrm{meas}}$, for the polar (spin-nematic squeezed) state. Each circle is centered at ($\langle Q_{yz}^{\mathrm{meas}}\rangle/N$, $\langle Q_{xz}^{\mathrm{meas}}\rangle/N$), with a radius of $\sqrt{3}\sqrt{(\Delta Q_{yz}^{\mathrm{meas}})^{2}+(\Delta Q_{xz}^{\mathrm{meas}})^{2}}/N$. (c-f) The mean values and standard deviations of the measured $Q_{xz}$ and $Q_{yz}$ at variable $\alpha$ with fixed $\Theta$ for the polar state (c,e) and spin-nematic squeezed state (d,f) as the probes. Each point represents the average of 100 repeated experiments.
@@
The experimental system is a spinor atomic BEC in the $F=1$ ground hyperfine manifold. Under the single spatial mode approximation, it can be described by the Hamiltonian [36]
@@
with $a_{m_{F}}$ and $a_{m_{F}}^{\dagger}$ the creation and annihilation operators for the $\left|F=1,m_{F}=\{0,\pm 1\}\right\rangle$ component. This Hamiltonian contains an interaction part and a quadratic Zeeman part. The first part includes a spin-mixing term generating paired atoms into $\left|F=1,m_{F}=\pm 1\right\rangle$ from $\left|F=1,m_{F}={0}\right\rangle$ and vice versa at a rate of $\left|c_{2}\right|$, and an energy shift term caused by elastic collisions. The second part is proportional to the atomic number operator $N_{0}$ and the coefficient $q$ is tunable by a bias magnetic field or an off-resonant MW dressing field [37].
@@
For a condensate prepared initially in the $\left|F=1,m_{F}=0\right\rangle$ state or the polar state, nearly all atoms remain in the $\left|F=1,m_{F}=0\right\rangle$ state during the short-term spin-mixing dynamics. In this undepleted pump regime, the total number of transferred atoms in $\left|F=1,m_{F}=\pm 1\right\rangle$ is small, the operator $a_{0}$ ($a_{0}^{\dagger}$) can be replaced approximately by a complex number $\sqrt{N}e^{-i\phi}$ ($\sqrt{N}e^{i\phi}$), and the above Hamiltonian reduces to the form of parametric down-conversion
$H=c_{2}(e^{-2i\phi}a_{1}^{{\dagger}}a_{-1}^{{\dagger}}+\mathrm{h.c.}),$ (2)
when $q$ is adjusted to compensate for the energy shift of the elastic collision at $q=\left|c_{2}\right|$. The interaction Eq. (2) produces a spin-nematic squeezed state as shown in Fig. 1(a), analog to the two-mode squeezed vacuum state [30, 31] in the widely discussed SU(1,1) interferometry. The sign of Eq. (2) can be easily flipped by inducing a $\pi$ phase shift of the $\left|F=1,m_{F}=0\right\rangle$ component to realize time-reversed evolution [38]. In this work, we demonstrate that the spin-nematic squeezed state, the atomic analog of the two-mode squeezed vacuum state, is sensitive to the spin rotation transformation $e^{-i(\varphi_{1}L_{x}+\varphi_{2}L_{y})}$ parameterized by two phases $\varphi_{1}$ and $\varphi_{2}$.
A complete description of the spin-1 system is provided by the SU(3) Cartesian dipole–quadrupole basis, which consists of the three components of the spin operators $L_{i}$ and the moments of the rank-2 quadrupoles or nematic tensors $Q_{ij}=L_{i}L_{j}+L_{j}L_{i}-(4/3)\delta_{ij}$, with $\{i,j\}\in\{x,y,z\}$ and $\delta_{ij}$ the Kronecker delta notation. In the undepleted pump regime, $N_{0}\sim N$, the collective spin operator $L_{x}$ ($L_{y}$) and the nematic operator $Q_{yz}$ ($Q_{xz}$) reduce to $e^{-i\phi}\sqrt{N}a_{S}^{\dagger}+\mathrm{h.c.}$ ($-ie^{-i\phi}\sqrt{N}a_{A}^{\dagger}+\mathrm{h.c.}$) and $-ie^{-i\phi}\sqrt{N}a_{S}^{\dagger}+\mathrm{h.c.}$ ($e^{-i\phi}\sqrt{N}a_{A}^{\dagger}+\mathrm{h.c.}$) approximately, which are proportional to the quadrature operators of the symmetric and asymmetric superpositions $a_{S}^{\dagger}=(a_{1}^{\dagger}+a_{-1}^{\dagger})/\sqrt{2}$ and $a_{A}^{\dagger}=(a_{1}^{\dagger}-a_{-1}^{\dagger})/\sqrt{2}$ of bosonic operators $a_{1}^{\dagger}$ and $a_{-1}^{\dagger}$, respectively.
@@
To read out the parameters $\varphi_{1}$ and $\varphi_{2}$, $Q_{yz}$ and $Q_{xz}$ need to be measured simultaneously, i.e., within one experimental run. For the measurement of quadrature operators of electromagnetic fields, homodyne detection represents a successful approach and is implemented by mixing the signal field to the coherent pump field [39]. Here in the spin-1 BEC system, atoms in $\left|F=1,m_{F}=0\right\rangle$ can be treated as a coherent pump field reference in measurements of the equivalent quadrature operators $Q_{yz}$ and $Q_{xz}$ [30, 40]. To realize such simultaneous read-out, $\pi/2$ coupling is applied between $a_{1}$ and $a_{-1}$ to form $a_{S}$ and $a_{A}$. Additionally, atoms in $|F=1,m_{F}=0\rangle$ are splitted into two coherent fields, through a tailored sequence of MW pulses with two orthogonal polarizations, which facilitate the coupling between $F=1$ manifold and the ancillary states within the $F=2$ manifold. This sequence leads to the transformations $a_{1,1}\rightarrow a_{S}$, $a_{1,-1}\rightarrow a_{A}$, and $a_{2,\pm 1}\rightarrow a_{0}/\sqrt{2}$. To avoid possible confusion with the state index in the above, the full label $\{F,m_{F}\}$ for the component mode is adopted here as opposed to the shorthand indexing by $m_{F}$ alone in the $F=1$ manifold.
The above seemingly cumbersome MW sequence is due to the limitations of our experimental setup, as we were unable to generate a microwave signal that was purely $\sigma_{+}$ or $\sigma_{-}$ polarized. Consequently, the $\pi/2$ coupling between $a_{1,1}$ and $a_{1,-1}$, as well as the redistribution of the atoms in $\left|F=1,m_{F}=0\right\rangle$ to the ancillary $|F=2,m_{F}=\pm 1\rangle$ is achieved through the application of the first three MW pulses shown in Fig. 1(c) by adjusting the power ratio between $\sigma_{+}$ and $\sigma_{-}$ polarizations. These transformations are to be implemented with high fidelity, the relative strength between $\sigma_{-}$ and $\sigma_{+}$ polarizations needs to be well controlled. The next two steps mix $a_{S}$ and $a_{A}$ to the pump mode and transfer the mode quadratures to the atomic population difference $N_{2,\pm 2}-N_{2,\pm 1}$. In the end, the read-out phase of the quadrature operators is determined by the phases of the second pulse and the pulses mixing $a_{S}$ ($a_{A}$) to the pump mode.
@@
Our experiment starts with a BEC containing $N=26500\pm 250$ ${}^{87}$b atoms in the $|F=1,m_{F}=0\rangle$ state under a bias magnetic field of 0.828 G and confined in a crossed dipole trap at $c_{2}=-2\pi\times 3.8$ Hz. The effective QZS $q=q_{B}+q_{\mathrm{MW}}$ includes contributions from both the bias magnetic field ($q_{B}$) and a blue-detuned microwave dressing field ($q_{\mathrm{MW}}$) which offers fast and precise control of $q$. The spin-mixing dynamics are initiated by quenching $q$ from $q\sim 13|c_{2}|$ to $q=|c_{2}|$ through adjusting the power of the dressing microwave. After a short time evolution, a spin-nematic squeezed state is generated and is followed by a $\pi/4$ spinor phase rotation around the $\left(Q_{yy}-Q_{zz}\right)/2$ and $\left(Q_{xx}-Q_{zz}\right)/2$ axis simultaneously, effected by two resonant microwave pulses between $|1,0\rangle$ and $|2,0\rangle$ with $\pi/4$ phase difference. The quadrature angle of optimal squeezing is then aligned to be most sensitively detected by $Q_{yz}$ and $Q_{xz}$, ensuring the probe state is most sensitive to the two-phase rotation around $L_{x}$ and $L_{y}$ axes.
@@
As illustrated in Fig. 1(b), spin-nematic squeezing is observed to build up gradually during spin-mixing evolution and the degree of squeezing can be quantified by the parameter

### v2

@@ Abstract
Quantum metrology employs entanglement to enhance measurement precision [1, 2, 3]. The focus and progress so far have primarily centered on estimating a single parameter. In diverse application scenarios, estimation of more than one single parameter is often required. Joint estimation of multiple parameters can benefit from additional advantages for further enhanced precision. Here we report quantum-enhanced estimation of simultaneous spin rotations around two orthogonal axes, making use of spin-nematic squeezing in an atomic Bose-Einstein condensate. Aided by the $F=2$ atomic ground hyperfine manifold coupled to the nematic-squeezed $F=1$ states as an auxiliary field through a sequence of microwave (MW) pulses, multiple spin-1 observables are simultaneously measured, reaching an enhancement factor 3.3 to 6.3 decibels (dB) beyond the classical limit over a wide range of rotation angles. Our work realizes the first quantum enhanced multiparameter estimation using entangled massive particles. The techniques developed and the protocols implemented also highlight the application of two-mode squeezed vacuum states in quantum-enhanced sensing of noncommuting spin rotations simultaneously.
@@
Quantum-enhanced sensing improves parameter estimation precision beyond the standard quantum limit (SQL), or the classical limit [1, 2, 3]. The fruitful recent achievement at advanced LIGO with squeezed light [4], yielding an increased number of detection events [5], has made quantum-enhanced sensing one of the most promising quantum information technologies in the near future. Such enhanced sensing has been demonstrated in experiments with photons [6], neutral atoms [7, 8, 9, 10, 11, 12, 13, 14, 15, 16], trapped ions [17, 18], and artificial atoms [19]. Significant technological gains are expected with atomic sensors such as in clocks [20] and magnetometers [21].
@@
Ideas for quantum-enhanced metrology have been extended to multiparameter settings as well, where estimation of multiple parameters can reach beyond SQL simultaneously [22, 23]. They are relevant for a broad range of applications, such as imaging and microscopy [24, 25], multidimensional field sensing [26, 27], and quantum networks for atomic clocks [28] etc., and they have motivated growing interests from both theoretical and experimental perspectives. A significant challenge in multiparameter estimation arises when the generators of parameter encoding do not commute. Observables that provide maximal information about different parameters need not necessarily commute [29], which could potentially impair quantum-enhanced sensitivities for all the parameters. Recently, Lorcán O. Conlon et al. have experimentally demonstrated optimal single- and two-copy collective measurements for simultaneously estimating two noncommuting qubit rotations [30].
Multimode squeezed states, which are often related to quantum entanglement and Einstein-Podolsky-Rosen steering, are essential resources for quantum-enhanced multiparameter estimation. The idea was first delineated in the form of quantum dense coding [31] and was later demonstrated experimentally with the two-mode squeezed vacuum state generated by optical parametric amplifiers [32, 33, 34, 35]. A related recent study [36] reports spatial displacement estimation beyond the SQL along two quadratures of its center-of-mass motion coordinate for a trapped ion. A two-mode squeezed state was recently prepared for a trapped ion motion using additional radio potentials without laser field and demonstrated SU(1,1) interferometry [37].
@@
Figure 1: Spin-nematic squeezing in both SU(2) subspaces. (a) The generation of spin-nematic squeezed state is illustrated by quasiprobability distributions on Bloch spheres for both SU(2) subspaces $\left\{L_{y},Q_{xz},\left(Q_{xx}-Q_{zz}\right)/2\right\}$ and $\left\{Q_{yz},L_{x},\left(Q_{yy}-Q_{zz}\right)/2\right\}$, simulated here for ${N=100}$ atoms. The initial polar state corresponds to all atoms in the $m_{F}=0$ component. Evolving at $q\approx|c_{2}|$, spin-nematic squeezing quickly develops along the separatrix (dark gray lines) on both Bloch spheres. (b) Schematics illustrating the simultaneous measurement of $Q_{xz}$ and $Q_{yz}$. From left to right, the first gray dotted line box illustrates the atomic number distribution of the initial nematic-squeezed state. Well-designed MW pulses realize a $\pi/2$-pulse coupling between $a_{1,+1}$ and $a_{1,-1}$ while coherently splitting atoms from $a_{1,0}$ into $a_{2,+1}$ and $a_{2,-1}$ effectively (enclosed in a red dashed box). The complete experimental MW sequence is detailed in SM. A small number of atoms remain in the $|F=1,m_{F}=0\rangle$ state due to imperfections in the pulse sequence. The next two gray dotted line boxes show the described $\pi/2$-pulse couplings between $|F=1,m_{F}\pm 1\rangle$ and $|F=2,m_{F}\pm 1\rangle$, as well as the transfer of $|F=1,m_{F}=\pm 1\rangle$ atoms to $|F=2,m_{F}=\pm 2\rangle$. (c) The measured squeezing parameter at the optimal squeezing (for $q=|c_{2}|$) versus evolution time. The gray dots show the experimental results in a single SU(2) subspace by applying an RF rotation to couple sublevels in the $F=1$ manifold [41]. The blue and orange dots show the experimental results of simultaneously measured squeezing in both SU(2) subspaces. The simultaneously observed squeezing parameter diminishes at longer times due to imperfect MW pulses as well as other decoherence or dissipation mechanisms. All markers denote experimental results averaged over 100 runs and the corresponding solid and dashed lines denote numerical calculations based on truncated Wigner approximation. The blue and orange dashed lines account for Gaussian-distributed noise in the amplitude of signal mode ($a_{S/A}$) and a Gaussian-distributed random phase in the pump mode ($a_{2,\pm 1}$), both calibrated as detailed in SM [38].
@@
Here we report quantum-enhanced simultaneous estimation of two unknown phases encoded in rotations about two orthogonal collective spin components $L_{x}$ and $L_{y}$ of a spin-1 atomic Bose-Einstein condensate (BEC). A complete description of the spin-1 system is provided by the SU(3) Cartesian dipole-quadrupole basis, which consists of the three components of the collective spin operators $L_{i}$ and the moments of the rank-2 quadrupoles or nematic tensors $Q_{ij}$ with $\{i,j\}\in\{x,y,z\}$. Detailed definitions of these operators are provided in Supplemental Material (SM) [38]. Through spin-mixing dynamics, an atomic analog of the two-mode squeezed vacuum state-a spin-nematic squeezed state-is generated [42, 41, 43], which exhibits squeezed quantum noise in both spin-nematic SU(2) subspaces spanned by $\left\{L_{y},Q_{xz},\left(Q_{xx}-Q_{zz}\right)/2\right\}$ and $\left\{Q_{yz},L_{x},\left(Q_{yy}-Q_{zz}\right)/2\right\}$. To estimate the phases encoded into the probe state, we implement a protocol to measure two quadrupole moments $Q_{xz}$ and $Q_{yz}$ simultaneously, reminiscent of atomic homodyne detection demonstrated earlier [42]. $Q_{yz}$ and $Q_{xz}$ can be simultaneously measured because $\langle L_{z}\rangle=0$ for the probe state used, which causes the two nematic operators to effectively commute, thus circumventing the limitations imposed by the Heisenberg uncertainty principle for noncommuting observables. By taking the $F=2$ hyperfine states as the auxiliary ones and applying a sequence of MW pulses to transform $Q_{xz}$ and $Q_{yz}$ into the population differences of atomic hyperfine sublevels while preserving the squeezed noise correlation in the protocol to be detailed later, both $Q_{xz}$ and $Q_{yz}$ are measured from a single absorption image of a BEC, and the metrological gains observed for joint estimation of the two phases reach 3.33.33.3 to 6.36.36.3 dB beyond the SQL over a wide range of rotations explored, where the SQL $1/\sqrt{2N}$ is for estimating two parameters with an ensemble of $N$ spin-1 particles [44] as detailed in SM. Consistent with the standard practice of estimating a single parameter beyond the classical limit, we carry out strong (collapsing) measurements to obtain the maximal amount of information by adopting atomic homodyne detection.
@@
Figure 2: Joint estimation of spin rotation angles around $\hat{L}_{x}$ and $\hat{L}_{y}$. (a) Schematics of the joint estimation interferometry protocol with quasiprobability distributions for the states plotted in the $\left\{L_{x},Q_{yz}\right\}$ and $\left\{L_{y},Q_{xz}\right\}$ planes, respectively. $U_{\mathrm{S}}$, $U_{\mathrm{R}}$, and $U_{\vec{\varphi}}$ denote the unitary evolutions generated by spin-mixing, spinor phase rotation, and encoding spin rotation, respectively. The measurement part is shown Fig. 1. (b) The distributions of the simultaneously measured data for the polar state (red dots) and spin-nematic squeezed state (blue dots) sampled over 100 times after spin rotations at different $\varphi_{1}$ and $\varphi_{2}$ encoded. The distribution of polar state is much broader than that of spin-nematic squeezed state, hence the higher sensitivities for the estimation from the latter. Here $\varphi_{1}=\alpha\cos\Theta$ and $\varphi_{2}=\alpha\sin\Theta$. The red (blue) circles denote the boundaries of the measured data, $Q_{yz}^{\mathrm{meas}}$ and $Q_{xz}^{\mathrm{meas}}$, for polar (spin-nematic squeezed) state. Each circle is centered at ($\langle Q_{yz}^{\mathrm{meas}}\rangle/N$, $\langle Q_{xz}^{\mathrm{meas}}\rangle/N$), with a radius of $\sqrt{3}\sqrt{(\Delta Q_{yz}^{\mathrm{meas}})^{2}+(\Delta Q_{xz}^{\mathrm{meas}})^{2}}/N$. (c)-(f) The mean values and standard deviations of the measured $Q_{xz}$ and $Q_{yz}$ at variable $\alpha$ and fixed $\Theta$ from using polar state (c,e) and spin-nematic squeezed state (d,f) as the probes. Each point represents 100 repeated experiment average.
@@
The system we consider is a spinor atomic BEC in the $F=1$ ground hyperfine manifold. Under the single spatial mode approximation [45], it can be described by the Hamiltonian
@@
with $a_{m_{F}}$ and $a_{m_{F}}^{\dagger}$ the creation and annihilation operators for condensed atoms of the $\left|F=1,m_{F}=\{0,\pm 1\}\right\rangle$ component. This Hamiltonian contains an interaction part and a single atom part. The former includes a spin-mixing term generating paired atoms into $\left|F=1,m_{F}=\pm 1\right\rangle$ from $\left|F=1,m_{F}={0}\right\rangle$ and vice versa at a rate of $\left|c_{2}\right|$, and an energy shift term caused by elastic collisions. The second part, which is proportional to the atomic number $N_{0}$ of $|F=1,m_{F}=0\rangle$, represents an effective quadratic Zeeman shift (QZS). Its coefficient $q=q_{B}+q_{\mathrm{MW}}$ includes contributions from both the bias magnetic field ($q_{B}$) and a blue-detuned microwave (MW) dressing field ($q_{\mathrm{MW}}$) for fast and precise control [46].
@@
Prepared initially in the $\left|F=1,m_{F}=0\right\rangle$ state or the polar state, nearly all atoms remain in the same state during the short-term spin-mixing dynamics. In this undepleted pump regime, the total number of atoms in $\left|F=1,m_{F}=\pm 1\right\rangle$ is small, the operator $a_{0}$ ($a_{0}^{\dagger}$) can be replaced approximately by $\sqrt{N}$, and the Hamiltonian  (1) reduces to the form of parametric down-conversion
$H=c_{2}(a_{1}^{{\dagger}}a_{-1}^{{\dagger}}+\mathrm{h.c.}),$ (2)
when $q$ is adjusted to compensate for the energy shift of the elastic collision at $q=\left|c_{2}\right|$. The interaction in Eq. (2) produces a spin-nematic squeezed state as shown in Fig. 1(a), analogous to the two-mode squeezed vacuum state [42, 41, 47] in the SU(1,1) interferometry. This state exhibits enhanced sensitivity to spin rotations parameterized by the two phases $\varphi_{1}$ and $\varphi_{2}$, specifically $e^{-i(\varphi_{1}L_{x}+\varphi_{2}L_{y})}$. For small rotation phases, the off-diagonal elements of the quantum Fisher information matrix become negligible, while the diagonal elements reduce to $4\langle L_{x}^{2}\rangle$ and $4\langle L_{y}^{2}\rangle$. When the spin-nematic state is rotated to induce squeezing along $Q_{yz(xz)}$ and antisqueezing along $L_{x(y)}$, the values of $4\langle L_{x(y)}^{2}\rangle$ exceed those of an unentangled polar state, thereby facilitating sensitivities beyond classical to both phases.
To jointly estimate both parameters $\varphi_{1}$ and $\varphi_{2}$, $Q_{yz}$ and $Q_{xz}$ need to be measured simultaneously, i.e., within one snapshot of absorption imaging. In the undepleted pump regime, $N_{0}\sim N$, the collective spin operator $L_{x}$ ($L_{y}$) and the nematic operator $Q_{yz}$ ($Q_{xz}$) reduce to the quadrature operators $\sqrt{N}a_{S}^{\dagger}+\mathrm{h.c.}$ ($-i\sqrt{N}a_{A}^{\dagger}+\mathrm{h.c.}$) and $i\sqrt{N}a_{S}^{\dagger}+\mathrm{h.c.}$ ($\sqrt{N}a_{A}^{\dagger}+\mathrm{h.c.}$), approximately, here $a_{S}^{\dagger}=(a_{1}^{\dagger}+a_{-1}^{\dagger})/\sqrt{2}$ and $a_{A}^{\dagger}=(a_{1}^{\dagger}-a_{-1}^{\dagger})/\sqrt{2}$ are symmetric and asymmetric superpositions of the bosonic operators $a_{1}^{\dagger}$ and $a_{-1}^{\dagger}$. For the measurement of quadrature operators of electromagnetic fields, homodyne detection represents a successful approach and is implemented by mixing the signal field to the coherent pump field [48]. Analogously, the two-phase spin rotation $e^{-i(\varphi_{1}L_{x}+\varphi_{2}L_{y})}$ reduces to simultaneous displacements corresponding to $a_{S}$ and $a_{A}$. It is worth emphasizing that $Q_{yz}$ and $Q_{xz}$ are quadrature operators of the two orthogonal modes $a_{S}$ and $a_{A}$, respectively. They can be simultaneously measured beyond the classical limit, thus circumventing the Heisenberg uncertainty constraint for noncommuting observables. Therefore, our measurement scheme is different from the scenario of simulteneously measuring $L_{x}$ and $Q_{yz}$ reported in Ref. [49], which effectively measure two conjugate quadratures in a single bosonic field $a_{S}$ and necessitates splitting the atomic population of the squeezed mode $a_{S}$ with auxiliary vacuum states via MW couplings, a process that introduces binomial atom-number fluctuations and limits the detectable squeezing to 3 dB below the SQL. Simultaneous measurement of quantum squeezing in both $Q_{yz}$ and $Q_{xz}$ can approach the intrinsic variance of the spin-nematic squeezed state theoretically, thereby facilitating near-optimal sensitivities of $\varphi_{1}$ and $\varphi_{2}$. Detailed analysis with numerical calculations are provided in the SM [38].
@@
Here in our system, atoms in $\left|F=1,m_{F}=0\right\rangle$ can be treated as a coherent pump field reference in measurements of the analogous quadrature operators $Q_{yz}$ and $Q_{xz}$ [42, 49]. To realize simultaneous measurement, a $\pi/2$-pulse coupling is applied between $a_{1}$ and $a_{-1}$ to form $a_{S}$ and $a_{A}$. Additionally, atoms in $|F=1,m_{F}=0\rangle$ are split into two coherent fields, through a tailored sequence of MW pulses with two orthogonal polarizations, which couples $F=1$ manifold and the ancillary states within the $F=2$ manifold. This sequence leads to the transformations $a_{1,1}\rightarrow a_{S}$, $a_{1,-1}\rightarrow a_{A}$, and $a_{2,\pm 1}\rightarrow a_{0}/\sqrt{2}$ as shown in Fig. 1(b). To avoid misunderstanding the state index used before, the full label $\{F,m_{F}\}$ for the component mode is adopted here instead of the shorthand indexing by $m_{F}$ alone in the $F=1$ manifold. Our current setup cannot generate pure $\sigma_{+/-}$ polarized microwaves, three MW pulses with controlled $\sigma_{+/-}$ power ratios are needed to complete the above transform (see SM).
@@
Our experiment starts with a BEC inside a bias magnetic field of 0.828 G. $N=26500\pm 250$ 87Rb atoms in the $|F=1,m_{F}=0\rangle$ state are confined in a crossed dipole trap at $c_{2}=-2\pi\times 3.8$ Hz. The spin-mixing dynamics are initiated by quenching $q$ from $q\sim 13|c_{2}|$ to $q=|c_{2}|$ through tuning the dressing microwave power. After a short time evolution, spin-nematic squeezed state is generated and a $\pi/4$ spinor phase rotation around the $\left(Q_{yy}-Q_{zz}\right)/2$ and $\left(Q_{xx}-Q_{zz}\right)/2$ axis is effected by two resonant MW pulses between |1,0|1,0$\rangle$| 1 , 0 ⟩ and |2,0|2,0$\rangle$| 2 , 0 ⟩ with $\pi/4$ phase difference. The quadrature angle of optimal squeezing is then aligned to be most sensitively detected by $Q_{yz}$ and $Q_{xz}$, ensuring the probe state is most sensitive to the two-phase rotation.
@@
As illustrated in Fig. 1(c), spin-nematic squeezing gradually builds up from spin mixing and the degree of squeezing can be quantified by the parameter